# Iterating $\mathsf{PP}$-packages without Choice: A Cohen symmetric seed and a localization framework

### v3

@@ Abstract
The Partition Principle $\mathsf{PP}$ asserts that whenever there is a surjection $A\twoheadrightarrow B$, there is an injection $B\hookrightarrow A$. Russell conjectured in 1906 that $\mathsf{PP}$ is equivalent to the Axiom of Choice $\mathsf{AC}$; while $\mathsf{AC}\Rightarrow\mathsf{PP}$ is immediate, the converse has remained open. We show that $\mathsf{PP}$ does not imply $\mathsf{AC}$ by constructing a transitive model of $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
Starting from a Cohen symmetric model $\mathcal{N}$ of $\operatorname{Add}(\omega,\omega_{1})$ built with a countable-support symmetry filter, we fix parameters $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ and perform a class-length countable-support symmetric iteration. At successor stages we use orbit-symmetrized packages that split targeted surjections, yielding $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$, while preserving $\mathsf{DC}$ and ensuring that $A$ remains non-well-orderable. A diagonal-cancellation/diagonal-lift infrastructure supplies a proper $\omega_{1}$-complete normal filter at limit stages. Finally, Ryan–Smith localization shows that under $\mathsf{SVC}^{+}(T)$, $\mathsf{PP}$ is equivalent to $\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}$, so the final model satisfies $\mathsf{PP}$ but not $\mathsf{AC}$.
@@ 1 Introduction
Whether the Partition Principle implies the Axiom of Choice is sometimes called “the oldest open problem in set theory.” The Partition Principle ($\mathsf{PP}$) asserts that whenever $A$ surjects onto $B$, one has $|B|\leq|A|$. Russell conjectured in 1906 that $\mathsf{PP}$ is equivalent to $\mathsf{AC}$, but no proof was given [1]. The converse is immediate: choice provides sections for all surjections. The question of whether $\mathsf{PP}$ and $\mathsf{AC}$ are equivalent over $\mathsf{ZF}$ has remained open.
Main Result. We resolve this question in the negative by constructing a model of
$\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC},$
thereby establishing that the Partition Principle does not imply the Axiom of Choice.
@@
The theory $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$ is consistent relative to $\mathsf{ZF}$.
Strategy.
The construction proceeds in two stages.
Stage 1: Seed model. We build a Cohen symmetric model $\mathcal{N}$ over $\operatorname{Add}(\omega,\omega_{1})$ using countable supports, producing a model of $\mathsf{ZF}+\mathsf{DC}$ containing a non-well-orderable set $A=\{c_{\alpha}:\alpha<\omega_{1}\}$ of Cohen reals. The fixed set $S$ witnesses $\mathsf{SVC}(S)$, which will drive the localization argument.
Stage 2: Symmetric iteration. We perform a class-length countable-support symmetric iteration over $\mathcal{N}$. At each successor stage, we force sections for surjections $f:Y\twoheadrightarrow X$ with $X,Y\subseteq T:=\mathcal{P}(S)$, as well as right inverses witnessing $\mathsf{AC}_{\mathsf{WO}}$. The iteration preserves $\mathsf{DC}$ and the non-well-orderability of $A$.
Localization. The global Partition Principle follows from Ryan–Smith’s localization theorem [17], which shows that under $\mathsf{SVC}^{+}(T)$, the principle $\mathsf{PP}$ is equivalent to the conjunction of $\mathsf{PP}\!\restriction\!T$ (the restriction to surjections between subsets of $T$) and $\mathsf{AC}_{\mathsf{WO}}$.
@@
Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes the axiom commonly called choice for well–orderable sets, in the sense of Howard–Rubin [7, Form 40]. That is:
Every well–orderable set of nonempty sets admits a choice function.
Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds: if $\neq\varnothing, Y\varnothing\notin Y$, and $|Y|\leq|\lambda|$, then there exists a function $f$ with domain $Y$ such that $f(y)\in y$ for all $y\in Y$.
This axiom should be distinguished from the trivial principle asserting choice for families already equipped with specified well–orderings, which is provable in $\mathsf{ZF}$. The present form asserts the existence of a uniform choice function for arbitrary well–orderable families, without any fixed well–orderings given in advance.
As shown by Pincus (see [7]), $\mathsf{AC}_{\mathsf{WO}}$ is equivalent to the statement that $\aleph(X)=\aleph^{*}(X)$ for all sets $X$. This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used implicitly in Ryan–Smith’s localization theorem [17].
Technical innovations.
Two devices are essential to making the iteration work.
(1) Orbit-symmetrized packages. A naïve attempt to force a section for a single surjection $f$ fails because the action of $\operatorname{Sym}(\omega_{1})$ on conditions does not extend to an action on the forcing poset (automorphisms would need to map $\mathbb{Q}_{f}$ to $\mathbb{Q}_{\pi\cdot f}$, a different poset). We resolve this by forcing with the full orbit package
$\mathbb{Q}_{[f]}:=\prod^{\mathrm{cs}}_{g\in[f]}\mathbb{Q}_{g},$
the countable-support product over all $g$ in the $\operatorname{Sym}(\omega_{1})$-orbit of $f$. This ensures that every $\pi\in\operatorname{Sym}(\omega_{1})$ acts by permuting coordinates and transporting structure, yielding a genuine automorphism of $\mathbb{Q}_{[f]}$.
(2) Diagonal-cancellation symmetries. Orbit-symmetrization alone does not make the generic section $s_{f}$ hereditarily symmetric: the naïve action of $\pi\in\operatorname{Fix}(E)$ on conditions still moves the $f$-coordinate generic. We introduce diagonal-cancellation automorphisms $\widehat{\pi}^{\,D}$ that act as the identity on a protected set $D$ of coordinates (including $f$) while acting as $\pi$ elsewhere. The stage filter includes subgroups $\Delta^{\uparrow}(E,\{f\})$ generated by these diagonal lifts, ensuring that the generic section is stabilized by a filter subgroup and hence belongs to the symmetric model. Crucially, diagonal lifts still act as $\pi$ on the Cohen base (stage 0), so they swap Cohen reals $c_{\beta}\leftrightarrow c_{\gamma}$ for $\beta,\gamma\notin E$, preserving the $\neg\mathsf{AC}$ witness.
Conventions.
Throughout, $\mathsf{DC}$ denotes $\mathsf{DC}_{\omega}$, i.e., dependent choice for $\omega$-sequences. We write $\mathsf{PP}\!\restriction\!T$ for the restriction of $\mathsf{PP}$ to surjections $f:Y\twoheadrightarrow X$ with both $X,Y\subseteq T$.
External tools.
We rely on three external frameworks, imported as black boxes.
(1) Symmetric iteration machinery. For the countable-support symmetric iteration framework—including successor and limit constructions, filter generation, $\omega_{1}$-completeness at limits, and the hereditarily-symmetric calculus—we cite Gilson [5]. See Remark 2.1 for the specific API items we invoke.
(2) DC preservation. For preservation of Dependent Choice in symmetric extensions, we follow Karagila [10] and Karagila–Schilhan [12], building on classical work of Pincus [14]. The criterion we use is: $\omega_{1}$-closed forcing with $\omega_{1}$-complete filter preserves $\mathsf{DC}$.
(3) Localization of $P$. We use Ryan–Smith’s local reflection framework [17], which provides the equivalence $\mathsf{PP}\Longleftrightarrow\mathsf{PP}\!\restriction\!T+\mathsf{AC}_{\mathsf{WO}}$ under $\mathsf{SVC}^{+}(T)$. For background on $\mathsf{SVC}$ and Hartogs–Lindenbaum spectra, see [16, 2].
For general background on forcing and symmetric extensions we follow Jech [8] and Kunen [13].
Outline.
Section 3 constructs the seed model $\mathcal{N}$ and verifies $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$. Section 4 recalls the Ryan–Smith localization theorem and defines the target principles $\mathsf{PP}\!\restriction\!T$ and $\mathsf{AC}_{\mathsf{WO}}$. Subsection 4.1 and Subsection 4.2 define the package forcings $\mathbb{Q}_{f}$ and $\mathbb{R}_{f}$ for forcing sections and right inverses. Section 5 constructs the symmetric iteration, including the orbit-symmetrization and diagonal-cancellation machinery, and assembles the final model $\mathcal{M}\models\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\neg\mathsf{AC}$.
2 Preliminaries and black-box iteration technology
We use the general symmetric-iteration machinery from [5] (and ultimately from Karagila’s iteration framework [9]). Here we only record the “API” we will use later. We treat the following as black boxes; the precise hypotheses and proofs are available in [5].
Remark 2.1 (Iteration API to be used later).
We will use, as needed, the following results from [5].
@@ 2 Preliminaries and black-box iteration technology
    Set-length and class-length versions of symmetric iterations (cf. [5, §2.1]).
@@
    Definitions and basic properties of countable-support symmetric iterations (cf. [5, Def. 3.1]).
@@
    A $\mathsf{DC}$-preservation theorem for the relevant support (cf. [5, Thm. 3.20]).
@@
    Generation and completeness/normality of limit filters in the relevant cofinality cases (cf. [5, Thm. 5.5]).
  5. (e)
    A characterization of HS names by “small support” assuming the stage model satisfies $\mathsf{DC}$ (cf. [5, Lem. 3.14]).
In our application, every stage model satisfies $\mathsf{DC}$ (Proposition 5.65), so the support characterization may be used freely throughout the iteration.
We emphasize that at each stage $\lambda$ the iteration produces a single ambient group $\mathscr{G}_{\lambda}$ acting by automorphisms on the full initial-segment forcing $\mathbb{P}_{\lambda}$ (API item (b)), so the action continues to move the original Cohen coordinates; we are not iterating by repeatedly forcing over intermediate symmetric models in isolation.
Our main tool for reducing the global Partition Principle to a local instance is the "Small Violations of Choice" framework.
Definition 2.2 (Small Violations of Choice, [17]).
Let $S$ be a set.
  1. 1.
    We say that $\mathsf{SVC}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is a surjection $f\colon S\times\eta\twoheadrightarrow X$ (equivalently, $|X|\leq^{*}|S\times\eta|$).
  2. 2.
    We say that $\mathsf{SVC}^{+}(S)$ holds if for every set $X$, there exists an ordinal $\eta$ such that there is an injection $f\colon X\hookrightarrow S\times\eta$ (equivalently, $|X|\leq|S\times\eta|$).
Ryan–Smith localization ingredients (used later). We will repeatedly appeal to two results from [17], which we record in Section 4 in the exact forms used in this paper:
  • Injective upgrade: for every set $S$, $\mathsf{SVC}(S)$ implies $\mathsf{SVC}^{+}(\mathcal{P}(S))$ (Fact 4.3).
  • Localization of $\mathsf{PP}$: assuming $\mathsf{SVC}^{+}(T)$, one has
    $\mathsf{PP}\ \Longleftrightarrow\ \bigl(\mathsf{PP}\!\restriction T\wedge\mathsf{AC}_{\mathsf{WO}}\bigr)$
    (Theorem 4.5, i.e. [17, Proposition 3.17]).
Accordingly, our strategy is to start from a Cohen symmetric seed $S$ with $\mathsf{SVC}(S)$, take $T:=\mathcal{P}(S)$, force $\mathsf{PP}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ while preserving $\mathsf{SVC}(S)$, and then invoke Theorem 4.5 to conclude global $\mathsf{PP}$.
@@ 3 The Cohen symmetric seed model
$\dot{c}_{\alpha}:=\{\langle\check{n},p\rangle:p(\alpha,n)=1\},c_{\alpha}=\dot{c}^{G}_{\alpha}$
the standard forcing to add $\omega_{1}$-many Cohen reals, presented as finite partial functions $p:\omega_{1}\times\omega\to 2$.
@@
$D_{\alpha,\beta}=\{p\in\operatorname{Add}(\omega,\omega_{1}):\exists n\,(p(\alpha,n)\neq p(\beta,n))\}.$
Given any $p$, choose $n$ such that neither $(\alpha,n)$ nor $(\beta,n)$ is in $\operatorname{dom}(p)$, and extend $p$ to $q\leq p$ with $q(\alpha,n)=0$ and $q(\beta,n)=1$. Then $q\in D_{\alpha,\beta}$. Hence $D_{\alpha,\beta}$ is dense, so any generic $G$ meets it and therefore $c_{\alpha}\neq c_{\beta}$. ∎

### v4

# A countable-support symmetric iteration separating $\mathsf{PP}$ from $\mathsf{AC}$

@@ Abstract
We construct, from a ground model of $\mathsf{ZFC}$, a transitive symmetric model $\mathcal{M}$ satisfying $\mathsf{ZF}+\mathsf{DC}+\mathsf{PP}+\mathsf{AC}_{\mathsf{WO}}+\neg\mathsf{AC}$. The construction starts with a Cohen symmetric seed model $\mathcal{N}$ over $\operatorname{Add}(\omega,\omega_{1})$ and performs an $\mathrm{Ord}$-length countable-support symmetric iteration. For fixed parameters $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ (as computed in $\mathcal{N}$), successor stages add orbit-symmetrized packages which force the localized splitting principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ (hence $\mathsf{PP}\restriction T$) and the choice principle $\mathsf{AC}_{\mathsf{WO}}$, while preserving $\mathsf{DC}$ and keeping $A$ non-well-orderable. A diagonal-lift/diagonal-cancellation scheme produces $\omega_{1}$-complete normal limit filters. A persistence argument yields $\mathsf{SVC}^{+}(T)$ in $\mathcal{M}$, and Ryan–Smith localization then upgrades $\mathsf{PP}\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ to $\mathsf{PP}$.
@@ 1 Introduction
The Partition Principle $\mathsf{PP}$ asserts that whenever there is a surjection $A\twoheadrightarrow B$, there is an injection $B\hookrightarrow A$. Over $\mathsf{ZFC}$ this is immediate, since every surjection admits a right inverse. The aim of this manuscript is to separate $\mathsf{PP}$ from the full Axiom of Choice by a symmetric-iteration construction.
Main theorem.
The following is proved as Theorem 5.73.
@@
Assume $V\models\mathsf{ZFC}$. In a class forcing extension $V[G]$ obtained by an $\mathrm{Ord}$-length countable-support symmetric iteration over a Cohen symmetric seed model $\mathcal{N}$, the final symmetric model $\mathcal{M}$ satisfies
$\mathsf{ZF}\ +\ \mathsf{DC}\ +\ \mathsf{PP}\ +\ \mathsf{AC}_{\mathsf{WO}}\ +\ \neg\mathsf{AC}.$
Metatheory.
All forcing and symmetry constructions are carried out over a fixed ground universe $V$. Set-sized stages are treated in the usual set-theoretic background. To define and analyze the $\mathrm{Ord}$-length iteration as a definable class forcing (and to use the forcing theorem for $\mathbb{P}_{\mathrm{Ord}}$), we work in a background class theory supporting the relevant recursions, namely $\mathsf{GBC}+\mathsf{ETR}$, using a class well-order of $V$ only where it is explicitly invoked for bookkeeping; see Section 2 and Remark 5.59.
@@
Throughout this paper, $\mathsf{AC}_{\mathsf{WO}}$ denotes choice for well-orderable sets: every well-orderable set of nonempty sets admits a choice function. Equivalently, for every ordinal $\lambda$, the principle $AC_{\lambda}$ holds. This is the form of $\mathsf{AC}_{\mathsf{WO}}$ used when applying Ryan–Smith localization (Theorem 4.5).
How $\mathsf{PP}$ is obtained.
Let $\mathcal{N}$ be the Cohen symmetric seed model over $\operatorname{Add}(\omega,\omega_{1})$ constructed in Section 3, and let $A=\{c_{\alpha}:\alpha<\omega_{1}\}$ be the canonical set of Cohen reals in $\mathcal{N}$. Fix $S:=A^{\omega}$ and $T:=\mathcal{P}(S)$ as computed in $\mathcal{N}$ (Definition 3.27); these parameters are held fixed throughout the iteration. Successor stages of the iteration schedule orbit-symmetrized package forcings which (in the resulting stage models) add:
  • witnesses for the localized splitting principle $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ (Definition 4.8); and
  • right inverses witnessing $\mathsf{AC}_{\mathsf{WO}}$ for well-orderable targets.
We force splitting because it is the local property that admits a uniform, highly closed "package forcing" $\mathbb{Q}_{f}$ per surjection, and it immediately implies the $\mathsf{PP}\restriction T$ hypothesis needed.
At limit stages, diagonal-lift/diagonal-cancellation symmetries are used to define $\omega_{1}$-complete normal limit filters compatible with countable support.
In the final model $\mathcal{M}$ we prove the local splitting statement (Proposition 5.65) and $\mathsf{AC}_{\mathsf{WO}}$ (Proposition 5.72), and we show that $\mathsf{SVC}(S)$ persists through the iteration (Lemma 5.69 and Corollary 5.70), hence $\mathsf{SVC}^{+}(T)$ holds in $\mathcal{M}$. Therefore, Ryan–Smith localization (Theorem 4.5) yields $\mathsf{PP}$ in $\mathcal{M}$ from $\mathsf{PP}\restriction T$ together with $\mathsf{AC}_{\mathsf{WO}}$; we obtain $\mathsf{PP}\restriction T$ because we force $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ (Definition 4.8). The remaining clauses of the main theorem are established by the corresponding final-stage propositions: $\mathsf{ZF}$ (Theorem 5.58), $\mathsf{DC}$ (Proposition 5.68), and $\neg\mathsf{AC}$ via non-well-orderability of $A$ (Proposition 5.67).
Consequences recorded here.
From $\mathsf{SVC}(S)$ and $\mathsf{PP}$ in $\mathcal{M}$, we derive the Ordering Principle (Corollary 5.74). We also record, conservatively, the Kinna–Wagner consequences proved in this manuscript (Remark 5.75).
Relation to countable-support symmetric iterations.
The iteration is an instance of the general countable-support symmetric-iteration framework developed in [5]; we use its successor and limit constructions, the $\omega_{1}$-completeness and normality facts for the generated filters, and the $\mathsf{ZF}/\mathsf{DC}$ preservation theorems. What is specific to the present manuscript is the package bookkeeping for $\mathsf{PP}^{\mathrm{split}}\!\restriction T$ and $\mathsf{AC}_{\mathsf{WO}}$ and the orbit-symmetrization required to obtain genuine automorphisms of the successor iterands (Remark 2.4).
What is not claimed.
Beyond the theory displayed above and the explicit corollaries proved in the final section (such as Corollary 5.74 and Remark 5.75), we do not claim additional choiceless consequences.
Organization.
Section 2 fixes the metatheoretic conventions and records the countable-support symmetric-iteration API items imported from [5]. Section 3 constructs the seed model $\mathcal{N}$ and establishes $\mathsf{DC}$, $\neg\mathsf{AC}$, and $\mathsf{SVC}(S)$ at stage 0. Section 4 recalls Ryan–Smith localization and isolates the fixed parameter $T$ together with the localized principles. Subsections 4.1 and 4.2 define the package forcings used at successor stages. Section 5 defines the class-length symmetric iteration and verifies the final-stage propositions, culminating in Theorem 5.73.
2 Preliminaries and metatheoretic conventions
We follow the forcing, automorphism, symmetric-extension, and countable-support symmetric-iteration framework developed in [5], based on Karagila’s iteration scheme [9]. This section fixes the metatheory and notation used throughout, and records precisely which results are imported from [5].
2.1 Background theory and conventions
Remark 2.1 (Metatheory vs. object theory).
All forcing notions, names, automorphism groups, filters, and supports are constructed in a fixed background universe $V$.
  • For the set-forcing and symmetric-extension statements used at bounded stages, we work over $V\models\mathsf{ZF}$.
  • Whenever we invoke preservation of dependent choice for a bounded stage symmetric model, we assume the background satisfies $V\models\mathsf{ZFC}$. Throughout this paper, $\mathsf{DC}$ means $\mathsf{DC}_{\omega}$ (dependent choice for $\omega$-sequences, in the standard set formulation used in [5, §3.5]).
  • For the class-length recursion through $\mathrm{Ord}$, we assume a background theory adequate for class recursion (e.g. $V\models\mathsf{GBC}+\mathsf{ETR}$), and we assume Global Choice only when explicitly invoked for bookkeeping.
When we say “countable”, $\omega_{1}$-complete”, etc., this is always with respect to the background universe $V$.
2.2 Restrictions, supports, and stage generics
Remark 2.2 (Forcing vs. group restrictions).
We follow the convention of [5, Remark 2.2] and do not overload restriction/projection notation.
  • For forcing, $p\upharpoonright\beta$ denotes the usual restriction of a condition $p\in\mathbb{P}_{\lambda}$ to an initial segment $\beta\leq\lambda$ (and when needed, $\pi_{\beta,\lambda}:\mathbb{P}_{\lambda}\to\mathbb{P}_{\beta}$ denotes the canonical projection map).
  • For groups, $\rho_{\beta,\lambda}:\mathscr{G}_{\lambda}\to\mathscr{G}_{\beta}$ denotes the restriction homomorphism.
Notation 2.3 (Stage generics).
Let $G\subseteq\mathbb{P}_{\mathrm{Ord}}$ be $V$-generic for the class-length iteration forcing. For each $\alpha\leq\mathrm{Ord}$, write $G_{\alpha}:=G\cap\mathbb{P}_{\alpha}$ for the induced $V$-generic filter on the initial segment forcing.
2.3 Imported iteration results from [5]
Remark 2.4 (Imported iteration facts).
We will use the following results from [5]. Each time we invoke an item below, we will explicitly point to the corresponding hypothesis in the current construction.
@@ 2.3 Imported iteration results from [5]
    Existence of set- and class-length iterations. The countable-support symmetric iteration recursion can be carried out through any set ordinal, and (under the background class-recursion assumptions used there) through $\mathrm{Ord}$; see [5, §Countable-support symmetric iterations and §Class-length iterations].
@@
    Framework and actions. At each stage $\lambda$ the iteration provides a single ambient group $\mathscr{G}_{\lambda}\leq\operatorname{Aut}(\mathbb{P}_{\lambda})$ acting by automorphisms on the full initial-segment forcing $\mathbb{P}_{\lambda}$ (so automorphisms may continue to move earlier “seed” coordinates); see [5, §Framework and Setup].
@@
    Limit filters. In the relevant cofinality cases, the canonical limit filter generated from earlier-stage filters is normal and $\omega_{1}$-complete; see [5, §Limit filters].
@@
    Bounded-stage $\mathsf{DC}$. For set ordinals $\lambda$, assuming $V\models\mathsf{ZFC}$, if the stage filter is $\omega_{1}$-complete then the stage symmetric model $M_{\lambda}=V(\mathbb{P}_{\lambda})_{\operatorname{HS}}$ satisfies $\mathsf{DC}$; see [5, §Dependent Choice and Choice Fragments].
We stress that [5] does not assert $\mathsf{DC}$ for the final class-length symmetric model in general. In this paper, $\mathsf{DC}$ for the final model is obtained by combining item (d) at bounded stages with the stage-localization lemma for set-sized $\mathbb{P}_{\mathrm{Ord}}$-names (Lemma 5.64).
Choice principles and the Ryan–Smith localization theorem are stated (and used) in Section 4.
@@ 3 The Cohen symmetric seed model
ordered by reverse inclusion.
For each $\alpha<\omega_{1}$, let $\dot{c}_{\alpha}$ be the canonical $\mathbb{P}$-name for the $\alpha$-th Cohen real,
$\dot{c}_{\alpha}\;:=\;\{\langle\check{n},p\rangle:p\in\mathbb{P}\ \wedge\ (\alpha,n)\in\operatorname{dom}(p)\ \wedge\ p(\alpha,n)=1\}.$
Fix $G_{0}\subseteq\mathbb{P}$ $V$-generic and write $c_{\alpha}:=\dot{c}_{\alpha}^{G_{0}}\in 2^{\omega}$.
@@
$D_{\alpha,\beta}=\Bigl\{p\in\mathbb{P}:\exists n\in\omega\ \bigl[(\alpha,n),(\beta,n)\in\operatorname{dom}(p)\ \wedge\ p(\alpha,n)\neq p(\beta,n)\bigr]\Bigr\}.$
Given any $p$, choose $n$ such that neither $(\alpha,n)$ nor $(\beta,n)$ is in $\operatorname{dom}(p)$, and extend $p$ to $q\leq p$ with $q(\alpha,n)=0$ and $q(\beta,n)=1$. Then $q\in D_{\alpha,\beta}$. Hence $D_{\alpha,\beta}$ is dense, so $G_{0}$ meets it and therefore $c_{\alpha}\neq c_{\beta}$. ∎